# End-to-End Breast Cancer Radiotherapy Planning via LMMs with Consistency Embedding

### v3

@@ IV-D Results
Figure 3: Qualitative comparison on 3D target volume segmentation task. Red arrows indicate errors.
TABLE VIII: Ablation study on adopting separate expertise for each textual task against unified strategy.
Training Strategy Internal (N=60) External (N=81)
R-1 / R-2 / R-L R-1 / R-2 / R-L
@@
Unified strategy 0.516 / 0.357 / 0.515 0.327 / 0.215 / 0.326
Distinct strategy 0.823 / 0.635 /0.822 0.805/ 0.634 / 0.804
Therapy Plan Suggestion
Unified strategy 0.510 / 0.265 / 0.502 0.409 / 0.197 / 0.406
Distinct strategy 0.652 / 0.491 / 0.649 0.606 / 0.450 / 0.606
TABLE IX: Ablation study on CESEG for target segmentation performance with input text variation.
@@
Dice $\uparrow$ IoU $\uparrow$
Ground Truth $0.860_{\pm 0.073}$ $0.761_{\pm 0.103}$
Generated $0.825_{\pm 0.101}$ $0.714_{\pm 0.134}$
Difference $\downarrow$ $0.035$ $0.047$
Ground Truth $0.847_{\pm 0.088}$ $0.743_{\pm 0.111}$
Generated ${0.840_{\pm 0.094}}$ ${0.733_{\pm 0.120}}$
Difference $\downarrow$ 0.007 0.010
V Ablation Study
Analysis on Strategy of Separate Expert To validate the rationale behind adopting separate strategies for each textual task, we conduct experiments by preparing another baseline model that unifies summarization and plan suggestion, and show the results in Table IX. We observe that the unified model’s performance is significantly inferior to our distinct training strategy. This implies that summarization task, which provides information within the input, is significantly different from suggesting treatment plan that does not exist within the input, indicating the challenging nature of the task. These results suggest that the knowledge required for each task necessitates task-aware training strategies.
@@ V Ablation Study
TABLE X: Inference computational complexity. RM: RO-LMM.
@@ VI Discussion
While our study provides valuable insights into the application of radiation oncology, several limitations must be acknowledged. Firstly, the dataset is confined to patients with initial breast cancer diagnoses. Expanding to diverse patient scenarios and other cancer types is crucial for broader applicability. Secondly, radiotherapy for breast cancer in real clinical setting often requires techniques like sequential boost or simultaneous integrated boost (SIB) to deliver higher radiation doses around the tumor bed. However, accurate tumor bed delineation requires integrating information from multiple imaging modalities, which simple plan-guided segmentation alone cannot achieve. Therefore, we did not include tumor bed delineation for boost techniques in this study to simplify the problem. Future research should focus on developing a framework capable of delineating these boost volumes to aid in clinical treatment. Finally, our model did not utilize all clinical information from the actual EMR for decision-making due to token length limitations. We only used MRI, US, and pathology results, considered most critical for initial diagnosis and treatment planning. Overcoming token length constraints to integrate all EMR data will be essential for future studies.
@@
Despite limitations, our RO-LMM demonstrates the potential to efficiently handle a broad range of clinical workflows using a single NVIDIA A6000 GPU within a 24GB memory, completing tasks in approximately 10 seconds (see Table X). This capability facilitates local implementation even with limited GPU resources and holds the promise of reducing the workload of radiation oncologists via the automation of clinical report-based plan suggestion for delineating radiation target volumes. Furthermore, our method demonstrates its effectiveness by surpassing closed-source LLMs on both a multi-center validation set and a publicly available dataset. These results suggest that our locally tuned RO-LMM-P++ holds significant potential to replace closed-source LLMs while maintaining data privacy and compliance with local data exposure regulations.
VII Conclusion
@@
In this work, we introduce RO-LMM, a multi-purpose, comprehensive foundation model tailored for radiation oncology. Addressing limitations in current medical AI models confined to specific tasks, RO-LMM demonstrates proficiency in diverse tasks encompassing overall workflow of radiation oncology: clinical report summarization, radiation treatment plan suggestion, and plan-guided 3D target volume segmentation. Another key contribution of this work is the introduction of consistency technique into both text and segmentation task. Results from multi-center cohort datasets confirm RO-LMM’s promising performance and noteworthy generalization capabilities across diverse tasks. These findings mark a significant stride toward developing a versatile AI model, hinting at the potential for a multi-purpose medical AI model in radiation oncology.
Ethical Statement All the patient information and any potential personally identifiable information from used datasets are de-identifiied, and clinical data usage is approved by the Institutional Review Board (IRB).
Appendix A Details of Training Dataset
@@
For data preparation, we gathered internal training and validation data from patients treated at Yonsei Cancer Center between January 2009 and December 2022. For external validation, data was collected from patients at Yongin Severance Hospital from January 2018 to December 2022, ensuring no overlap with the training set. For pre-processing CT scans, all the chest CT scans and target volumes were re-sampled with identical voxel spacing. The intensity values of the CT scans were truncated between -1,000 and 1,000 of Hounsfield unit (HU), and linearly normalized to the range of 0 to 1.0. During training, 3D patches was randomly cropped, whereas, during inference, the entire 3D CT scan was tested using sliding windows of the same spatial dimension as the training patches.
@@ Appendix C Detailed Prompt and Rubric for Evaluation
We design prompts for evaluating generated treatment plans using a reference-guided scoring rubric curated by clinical experts, as detailed in Table A1. Our approach builds on previous GPT-based evaluation methods and integrates clinical experts’ knowledge into the prompts. Using these tailored prompts, we assess the performance of our RO-LMM model and all baseline methods with GPT-4.0.
TABLE A1: Score rubrics for treatment plan suggestion.

### v4

@@ 5 Results
5.3 3D Target Volume Segmentation.
As reported in Table 5, we first compare the performance of various 3D segmentation models for target volume segmentation. Notably, both 3D SegMamba [52] and 3D UNETR [16] demonstrate inferior performance compared to the 3D Residual U-Net. Transformer-based or Mamba-based architectures significantly increase computational memory requirements, which constrains the input image dimensions. In the context of radiotherapy segmentation, where the network must process and align volumetric information in the entire volume-level, such input size limitations can be regarded as a significant limitation, especially under constrained computational resources. Consequently, we selected the more efficient 3D Residual U-Net as the backbone to maximize the 3D input dimensions and ensure optimal performance. Furthermore, our proposed plan-guided 3D target segmentation frameworks (RO-LMM-SEG/SEG+/SEG++) excel other default methods with large gains, specifically remained stable performance across the external validation setting. This can be also observed in qualitative results. As shown in Figure 3(a) and (c), despite the ground truth mask posing target volume on the one side of the breast, the unimodal 3D Residual U-Net incorrectly segments contours not only the target volume but also the outside of the target breast, while all the multimodal RO-LMM variants correctly segment target volume aligning consistently with the ground truth masks. Moreover, as shown in Figure 3(a) and (b), RO-LMM-SEG+ with NESEG module yields noisy segmentation output as indicated as red arrows, and RO-LMM-SEG sometimes under-segments the target volume. In contrast, our proposed RO-LMM-SEG++ with CESEG module accurately contours the breast and regional lymph nodes that need to be treated. Through experiments on consecutive tasks, including radiotherapy strategy suggestion and 3D target volume segmentation, we found that simple noise injection strategies (NEFT and NESEG) are not consistently effective across all tasks. In contrast, our proposed CEFT and CESEG methods demonstrate robustness and offer a more generalizable solution within the ROB-LMM framework.
We further categorize all patients into two distinct types: patients those who had breast-conserving surgery (typical type) and patients those who underwent total mastectomy (atypical type). The comparative results are presented in the Table 6 separately, with the unimodal model referred to the 3D Residual U-Net and the multimodal model denoted as our proposed RO-LMM-SEG++ with CESEG. Firstly, our proposed multimodal model shows substantial performance improvements of approximately up to 5% and 10% for internal and external validation, respectively. When comparing the typical and atypical types of patients, particularly in the less prevalent atypical type during external validation, we observed notable enhancements with a performance gain of up to 20%. This substantial generalizbility observed in extremely low-prevalence distributions demonstrates the data efficiency of our multimodal concept, leveraging textual information compared to the traditional unimodal approach. Additionally, we specifically categorize cases where regional lymph node treatment is not required and the CTV definition includes only the breast in Table 6, denoted as breast-only, and compare our results with other studies focused on breast segmentation for radiotherapy, as presented in [44]. Our method demonstrates comparable performance, achieving a Dice score of 0.891 and an HD-95 of 2.985 mm for the internal dataset.
5.4 Inference on Public Dataset.
To validate the effectiveness of our methods when AI-generated results are reused as input, we evaluated the performance by inputting each model’s report summarization results its plan suggestion task. We conducted experiments on a publicly available text-related dataset, as detailed in Table 7. Our proposed methods consistently outperformed other specialized medical LLMs and variants of GPT (such as ChatGPT and GPT-4) in both clinical report summarization and radiotherapy strategy suggestion tasks. Notably, our RO-LMM-P++ method demonstrated substantial performance enhancements, with improvements of up to 0.55 and 0.31 over other medical LLMs and GPT variants, respectively, demonstrating our model’s robustness to error accumulation during sequential generations.
6 Ablation Study
6.1 Analysis on Strategy of Separate Expert
To validate the rationale behind adopting separate strategies for each textual task, we conduct experiments by preparing another Default model that unifies summarization and plan suggestion, and show the results in Table 8. We observe that the unified model’s performance is significantly inferior to our distinct training strategy. This implies that summarization task, which provides information within the input, is significantly different from suggesting radiotherapy strategy that does not exist within the input, indicating the challenging nature of the task. These results suggest that the knowledge required for each task necessitates task-aware training strategies.
Table 8: Ablation study on adopting separate expertise for each textual task against unified strategy.
Training Strategy Internal External
R-1 $\uparrow$ R-2 $\uparrow$ R-L $\uparrow$ R-1 $\uparrow$ R-2 $\uparrow$ R-L $\uparrow$
@@ 6.1 Analysis on Strategy of Separate Expert
Unified strategy 0.516 0.357 0.515 0.327 0.215 0.326
Distinct strategy 0.783 0.605 0.783 0.788 0.607 0.788
Radiotherapy Strategy Suggestion
Unified strategy 0.510 0.265 0.502 0.409 0.197 0.406
Distinct strategy 0.655 0.500 0.655 0.615 0.459 0.615
Table 9: Ablation study on CESEG for target segmentation performance with input text variation.
@@
Dice $\uparrow$ (CI) surfaceDice $\uparrow$ (CI) IoU $\uparrow$ (CI)
Ground Truth 0.843 (0.823-0.863) 0.559 (0.527-0.591) 0.738 (0.711-0.764)
Generated 0.818 (0.795-0.841) 0.524 (0.493-0.555) 0.705 (0.676-0.736)
Difference $\downarrow$ $0.025$ $0.035$ $0.033$
Ground Truth 0.842 (0.819-0.861) 0.558 (0.524-0.589) 0.736 (0.708-0.761)
Generated 0.838 (0.818-0.858) 0.555 (0.523-0.583) 0.732 (0.706-0.757)
Difference $\downarrow$ 0.004 0.003 0.004
6.2 Analysis on Consistency Regularization
@@
6.3 Analysis of Noise type and Intensity ($\alpha$)
Since our proposed method involves injecting noise into embeddings, we conduct a study to analyze the effects of both noise type and intensity under various conditions, as shown in Table 10. First, regarding noise type, we observe that the method performed best with uniform noise compared to Gaussian noise, leading us to adopt uniform noise as the default configuration. We then examine noise intensity by varying the parameter $\alpha$ from 5 to 15. Lower values of $\alpha$ result in improved performance on both internal and external datasets, while excessively high values caused a noticeable drop in performance on the external dataset. Building on the main motivation behind NEFTune [20], we hypothesize that introducing noise to the embedding during training reduces overfitting to the specific characteristics of the instruction fine-tuning dataset, such as formatting details and exact wording. This allows the model to generate responses that incorporate broader knowledge rather than simply replicating the instruction distribution. However, increasing the noise intensity excessively can cause a mismatch between the original data distribution and the test set distribution, underscoring the importance of selecting an optimal noise intensity. Based on these empirical results, we selected $\alpha=5$ as the baseline.
6.4 Analysis of Finetuning Approach and Model Backbone
In our work, we adopt the full-finetuning approach for instruction finetuning of LLMs. To validate this choice, we conduct experiments comparing full-finetuning to parameter-efficient methods, such as LoRA as indicated in Table 10. The results showed that using LoRA led to a significant drop in performance on both internal and external datasets. Given the specialized nature of our scope, focused on medical professionals, parameter-efficient finetuning proved insufficient for capturing the target domain knowledge. These findings indicate that full-finetuning is essential for adapting LLMs to the medical domain, particularly in the context of radiation oncology for breast cancer.
We conduct experiments using the larger 13B LLaMA-2 backbone version and observe a similar performance trend compared to the 7B model as shown in Table 10. Despite the larger backbone size, our proposed CEFT method outperforms the other baselines, demonstrating scalability. Taking into account both cost-effectiveness and performance, we adopt the 7B model as the default configuration.
Table 10: Component analysis of our proposed method on radiotherapy strategy suggestion.
Component Condition Ours Internal Validation External Validation
R-1 $\uparrow$ R-2 $\uparrow$ R-L $\uparrow$ R-1 $\uparrow$ R-2 $\uparrow$ R-L $\uparrow$
Noise type Uniform 5 $\checkmark$ 0.655 0.500 0.655 0.615 0.459 0.615
Gaussian 5 0.625 0.449 0.629 0.588 0.425 0.586
Noise Intensity ($\alpha$) 5 $\checkmark$ 0.655 0.500 0.655 0.615 0.459 0.615
10 0.648 0.488 0.647 0.597 0.435 0.598
15 0.642 0.481 0.642 0.576 0.415 0.572
Finetuning Full-finetune $\checkmark$ 0.655 0.500 0.655 0.615 0.459 0.615
LoRA 0.384 0.352 0.373 0.325 0.301 0.323
Backbone LLaMA-13B 0.645 0.478 0.640 0.597 0.436 0.594
+NEFT 0.658 0.500 0.653 0.601 0.438 0.597
+CEFT $\checkmark$ 0.662 0.502 0.661 0.613 0.441 0.603
Table 11: Inference computational complexity.
@@ 7 Discussion
While our study provides valuable insights into the application of radiation oncology, several limitations must be acknowledged. Firstly, the dataset is confined to patients with initial breast cancer diagnoses. Expanding to diverse patient scenarios and other cancer types is crucial for broader applicability. Secondly, radiotherapy for breast cancer in real clinical setting often requires techniques like sequential boost or simultaneous integrated boost (SIB) to deliver higher radiation doses around the tumor bed. However, accurate tumor bed delineation requires integrating information from multiple imaging modalities, which simple plan-guided segmentation alone cannot achieve. Therefore, we did not include tumor bed delineation for boost techniques in this study to simplify the problem. Future research should focus on developing a framework capable of delineating these boost volumes to aid in clinical treatment. Thirdly, the inclusion of non-standard dose prescriptions reflecting country- or institution-specific practices in the training data can limit the generalizability of the results and cause confusion for non-experts. To ensure safer and more broadly applicable clinical implementation, it is crucial to train models using standardized, evidence-based prescriptions. Finally, our model did not utilize all clinical information from the actual EMR for decision-making due to token length limitations. We only used MRI, US, and pathology results, considered most critical for initial diagnosis and treatment planning. Overcoming token length constraints to integrate all EMR data will be essential for future studies.
@@
Despite limitations, our RO-LMM demonstrates the potential to efficiently handle a broad range of clinical workflows using a single NVIDIA A6000 GPU within a 24GB memory, completing tasks in approximately 10 seconds (see Table 11). This capability facilitates local implementation even with limited GPU resources and holds the promise of reducing the workload of radiation oncologists via the automation of clinical report-based plan suggestion for delineating radiation target volumes. Furthermore, our method demonstrates its effectiveness by surpassing closed-source LLMs on both a multi-center validation set and a publicly available dataset. These results suggest that our locally tuned RO-LMM-P++ holds significant potential to replace closed-source LLMs while maintaining data privacy and compliance with local data exposure regulations.
8 Conclusion
@@
In this work, we introduce RO-LMM, a multi-purpose, comprehensive foundation model tailored for radiation oncology. Addressing limitations in current medical AI models confined to specific tasks, RO-LMM demonstrates proficiency in diverse tasks encompassing overall workflow of radiation oncology: clinical report summarization, radiotherapy strategy suggestion, and plan-guided 3D target volume segmentation. Another key contribution of this work is the introduction of consistency technique into both text and segmentation task. Results from multi-center cohort datasets confirm RO-LMM’s promising performance and noteworthy generalization capabilities across diverse tasks. These findings mark a significant stride toward developing a versatile AI model, hinting at the potential for a multi-purpose medical AI model in radiation oncology.
Acknowledgments
This research was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (#2022R1A2C2008623), and also by the Ministry of Education (#RS-2023-00242164). This work was also supported by the Korea Health Industry Development Institute (KHIDI) grant funded by the Ministry of Health & Welfare, Republic of Korea (#HI23C0730000023). This work was also supported by the National Research Foundation of Korea (NRF) (#RS-2023-00262527) and the Korea Medical Device Development Fund grant funded by the Korea government (the Ministry of Science and ICT, the Ministry of Trade, Industry and Energy, the Ministry of Health & Welfare, the Ministry of Food and Drug Safety) (Project Number: 1711137899, KMDF_PR_20200901_0015).
Appendix A Details of Training and Validation Dataset
For data preparation, we gathered internal training and validation data from patients treated at Yonsei Cancer Center between January 2009 and December 2022.
@@ Appendix A Details of Training and Validation Dataset
For external validation, data was collected from patients at Yongin Severance Hospital from January 2018 to December 2022, ensuring no overlap with the training set. For pre-processing, all the chest CT scans and target volumes were re-sampled with identical voxel spacing. The Hounsfield unit (HU) of the CT scans were truncated between -1,000 and 1,000, and linearly normalized to the range of 0 to 1.0. During training, 3D patches was randomly cropped, whereas, during inference, the entire 3D CT scan was tested using sliding windows of the same spatial dimension as the training patches. All the patient information and any potential personally identifiable information from used datasets are de-identifiied, and clinical data usage is approved by the Institutional Review Board (IRB).
@@ Appendix C Detailed Prompt and Rubric for Evaluation
We design prompts for evaluating generated report summarizations and radiotherapy strategies using a reference-guided scoring rubric curated by clinical experts, as detailed in Table A1 and  A2, respectively. For the report summarizations, there are four evaluation criteria, with the first criterion having a weight of 2, making the total possible score 5. For the radiotherapy strategy suggestions, there are five criteria, each contributing a score of 1, resulting in a total score of 5. Our approach builds on previous GPT-based evaluation methods and integrates clinical experts’ knowledge into the prompts. Using these tailored prompts, we assess the performance of our RO-LMM model and all Default methods with GPT-4.0.
Table A1: The proposed expertise-based rubrics for assessing the performance of clinical report summarization.
Based on rubrics, score each clinical report summarization.
Rubric 1: Relevance Assessment of whether the necessary information from the Ground Truth (GT) is included, and identification of any extraneous information
Good (2): The information from the GT is appropriately included.
Partially correct (1): Some relevant information is included, but there are omissions or irrelevant details.
Mostly irrelevant (0): The majority of the information is unnecessary or irrelevant.
Rubric 2: Formatting Evaluation of whether the necessary information is presented in the correct format and location:
Incorrectly placed (0): Information is not presented in the appropriate format or location.
Correctly placed (1): The necessary information is accurately positioned in the expected format.
Rubric 3: Length Assessment of whether the writing is excessively verbose:
Too verbose (0): The writing is overly lengthy and lacks conciseness.
Sufficiently summarized (1): The content is effectively summarized without unnecessary verbosity.
Rubric 4: Consistency Identification of conflicting information between sources:
Present (0): Conflicting pieces of information are observed.
Absent (1): There are no conflicting pieces of information between sources.
Table A2: Score rubrics for radiotherapy strategy suggestion.
@@
[Abbreviations]
RT. Radiotherapy
BCS. Breast Conserving Surgery
WB. Whole Breast
CW. Chest Wall
AXL. Axilla Lymph Node
SCL. Supraclavicular Lymph Node
IMN. Internal Mammary Lymph node
RTOG. Radiation Therapy Oncology Group
ESTRO. European Socity for Radiotherapy and Oncology
SIB. Simultaneous Integrated Boost technique, same as tumor bed boost (sequential), fast-forward protocol
[Ground Truth Answer Begin]
$<$Ground Truth Plan$>$
[Ground Truth Answer End]
[Model Responses Begin]
$<$Model Generated Plan$>$
[Model Responses End]
[Answer Format]
Model, Answer for R1, Answer for R2, Answer for R3, Answer for R4, Answer
for R5, scores ONLY for each Rubric in a tabular format, without explanation: